# STORM: A Spatio-Temporal Factor Model Based on Dual Vector Quantized Variational Autoencoders for Financial Trading

### v3

@@ 5.3. Downstream Task Performance (for RQ2)
We use three types of metrics: profit, risk, and risk-adjusted profit to evaluate STORM’s investment performance on downstream tasks, i.e., portfolio management and algorithmic trading.
@@
For profitability, STORM outperforms all the baselines, with average improvements of 106.10% on APY over the best-performing baseline in the PM task and 43.64% over the best baseline methods on five stocks in the AT task. These results underscore the model’s capability to extract meaningful factors, and optimize decision-making for enhanced financial performance. For risk resistance, although STORM does not outperform the best baseline methods in terms of MDD and AVO on the INTC and MSFT datasets for the trading task, this can be attributed to the extreme price fluctuations during the COVID-19 event. Both stocks experienced a sharp decline of over 30%, followed by a subsequent rebound. Given STORM’s strong profitability capability, its trading strategy tends to aggressively capture trends, which naturally leads to larger drawdowns and increased volatility as a trade-off for higher potential returns.
@@
Moreover, considering the return and risk simultaneously, STORM performs the best on ASR and CR with an average improvement of 58.64% and 166.83% respectively, over the best baseline result in the PM task, and surpasses the best baseline results across three stocks in the AT task. Such an outstanding ability to balance return and risk is particularly valuable in real-world investment, where achieving high returns while effectively managing risk is crucial for long-term financial success.
@@ 5.4. Ablation Study (for RQ3)
We conducted ablation experiments with two simplified versions of STORM, namely STORM-w/o-TS and STORM-w/o-CS, which only extract cross-sectional and time-series factors, respectively. Specifically, the absence of either time-series or cross-sectional factors significantly reduces the model’s effectiveness across various tasks. In the stock return prediction task, STORM outperforms both simplified models, with an average improvement of 21.73% on the RankIC metric and 24.54% on the RankICIR metric, highlighting the enhanced accuracy achieved by incorporating both time-series and cross-sectional factors. In the PM and AT tasks, STORM and two simplified models all performed well. Notably, STORM-w/o-TS exhibited the lowest AVO among the three models, although its profitability did not surpass that of STORM. These findings suggest that methods like FactorVAE (Duan et al., 2022), which focus solely on cross-sectional factors, may struggle to capture the full complexity of the financial market. Therefore, integrating both time-series and cross-sectional factors is essential for robust and accurate financial modeling.
@@ 5.5. Discussion on Performance
Efficiency. STORM achieves remarkable efficiency in real-world trading scenarios. Unlike high-frequency trading systems requiring ultra-low latency and specialized hardware, our daily trading strategy operates efficiently on standard CPU hardware. This is enabled by a compact model architecture characterized by: i) a minimalist design with only a few transformer blocks (see Appendix B), and ii) a lightweight parameterization comprising 15.2 million FP32 parameters on DJ30. This design ensures seamless deployment on edge devices while supporting concurrent market analysis and trading.
Effectiveness. STORM exhibits exceptional performance compared to baseline methods, showcasing versatility across tasks including PM and AT. Unlike baselines that often excel in specific tasks but lack generalizability, STORM consistently delivers robust overall results. Moreover, STORM can balance returns and risks by emphasizing comprehensive investment performance over individual financial metrics. Collectively, these attributes substantiate that the additional computational cost incurred by STORM is well justified by its robustness and adaptability across diverse tasks and datasets.
Generalization. STORM demonstrates robust generalizability by achieving strong performance across markets with varying scales and distinct characteristics. It excels on both the DJ30, which reflects large-cap industrial blue chips, and the SP500, representing a broader market across diverse industries. Furthermore, its flexible encoder-decoder architecture allows the integration of additional layers such as Mamba  (Gu and Dao, 2023; Dao and Gu, 2024) and ensures applicability to diverse trading environments and asset classes.
@@ 6. Conclusions and Future Work
In this paper, we propose STORM, a spatio-temporal latent factor model that leverages the dual VQ-VAE framework to learn both time-series and cross-sectional factors. STORM ensures the orthogonality and diversity of factors, which are crucial for effective factor selection in financial trading. Extensive experiments demonstrate the superiority of STORM over the state-of-the-art methods across a prediction task and two downstream tasks, showcasing its robust performance and effectiveness. In the future, we plan to extend STORM by incorporating more side information, e.g., news, to explore the effects of exogenous factors.
Appendix A Notations
We provide the main notations in Table A.5 used in the paper.
Notation Description
$N$ total number of stocks
$W$ window size of the observation interval
$D$ total number of features of data
$T$ total number of trading days
$K$ codebook size
$p_{i,t}$ price of stock $i$ at day $t$
$\mathbf{x}$ observed data, which contains stock historical price data $\mathbf{p}$ and indicators $\mathbf{d}$
$\mathbf{x}^{\prime}_{ts}$, $\mathbf{x}^{\prime}_{cs}$, reconstructed observed data of the TS and CS modules
$\hat{\mathbf{y}}_{i,t}$, $\mathbf{y}_{i,t}$ predicted and true return of stock $i$ at day $t$
$\mathbf{w}_{t}$ the portfolio weight vector for $N$ stocks at trading day $t$
$a_{i,t}$ action of day $t$ on the stock $i$
$s_{t}$ state of day $t$
$\mathbf{z}_{e}^{ts}(\mathbf{x})$, $\mathbf{z}_{e}^{cs}(\mathbf{x})$ latent embeddings of the TS and CS modules
$\mathbf{z}_{q}^{ts}(\mathbf{x})$, $\mathbf{z}_{q}^{cs}(\mathbf{x})$ quantized embeddings of the TS and CS modules
$\mathbf{z}_{e}(\mathbf{x})$ factor embeddings
$\mathbf{z}$ posterior or latent factors
$k$ total number of selected stocks in PM strategy
$d$ the max number of stocks to be sold out
Table A.5. Notations in the paper.
Appendix B Implement Details
We implement all experiments by using PyTorch and conduct the experiments on two 80GB memory NVIDIA RTX H100 GPUs. The code is provided in the repository  [^4] https://anonymous.4open.science/r/Storm-2823. The input data is organized into batches with dimensions $(B,W,N,D)$, where $B=16$ denotes the batch size. $N$ represents the number of stocks, 28 for the DJ30 index and 408 for the SP500 index. The dimension $W=64$ corresponds to the number of historical data days, and $D=152$ represents the total number of features, including the OHLC data, technical indicators, and temporal information.
For the time-series module, we use an 4-day period for each stock as a patch, meaning the patch size is $(4,1,D)$. In contrast, for the cross-sectional module, we use $N$ stocks for each day as a patch, so the patch size is $(1,N,D)$. For both the encoders and decoders, we employ stacked transformer blocks. Each encoder consists of 4 layers of blocks with 4 attention heads, while each decoder consists of 2 layers of blocks with 8 attention heads. The size of both codebooks is selected from the set {128, 256, 512, 1024, 2048}, with 512 yielding the best results. A consistent embedding dimension of 256 is employed across the encoders, decoders, and codebooks.
For the optimizer, we use AdamW with a learning rate of $1\times 10^{-4}$ and a weight decay of 0.05. The learning rate starts at 0, ramps up to $1\times 10^{-4}$ over 100 epochs with a linear warmup scheduler, and then decays to 0 over 1000 epochs. For the loss function balancing, the coefficients for the clip loss, as well as the price reconstruction losses for both the TS and CS modules, are all set to $1\times 10^{-3}$. The commitment loss coefficient for the codebooks is 1.0, while the coefficients for the orthogonality loss and diversity loss are both set to 0.1. Additionally, the coefficient for the price prediction loss of the factor module, as well as the KL divergence loss between the posterior and prior distributions, is also set to 0.1.
For downstream tasks, we employ the TopK-Drop rule strategy for portfolio management, using the recommended parameters from Qlib, where $k=5$ and $d=3$. For the trading task, we adopt the Proximal Policy Optimization (PPO) RL strategy. The CS and TS factors from the pretraining phase are used as the state. Notably, we retain all CS factors, which represent market characteristics, while only preserving the TS factors relevant to the currently traded stocks, reflecting stock-specific characteristics. The action space has a dimension of 3, corresponding to the choices of buy, hold, and sell. Our reward calculation considers value to account for scenarios where remaining cash may not cover the purchase of a new stock. Relying solely on price changes without considering position and cash factors is unreasonable. Value is defined as $v=p\times m+c$ where $p$ is the current price, $m$ is the position, and $c$ is the remaining cash. The reward is calculated as $r=\frac{v_{\text{post}}-v_{\text{pre}}}{v_{\text{pre}}}.$ Unlike compound interest, our method measures the rate of change from the initial value to the current value. For both downstream tasks, we start with an initial cash of $1\times 10^{6}$, while also accounting for a transaction cost ratio of $1\times 10^{-4}$.
For the pretraining prediction tasks and the downstream portfolio management tasks, each set of experiments is conducted 10 times with different random seeds, and the mean and variance are reported. For the trading tasks, each set of experiments is conducted 5 times with different random seeds, and the mean of the evaluation metrics is reported. Notably, the default parameters provided by the official implementations are used for baseline models, ensuring a consistent and fair comparison across different approaches.

### v4

@@ 5.2. Performance on Investment Tasks (for RQ1)
We use three types of metrics: profit, risk, and risk-adjusted profit to evaluate STORM’s investment performance on downstream tasks, i.e., portfolio management and algorithmic trading. The evaluation results are presented in Table 2 and Table 3.
@@
For profitability, STORM outperforms all the baselines, with average improvements of 116.36% on APY over the best-performing baseline in the PM task and 43.64% over the best baseline methods on five stocks in the AT task. These results underscore the model’s capability to extract meaningful factors, and optimize decision-making for enhanced financial performance. For risk resistance, although STORM does not outperform the best baseline methods in terms of MDD and AVO on the INTC and MSFT datasets for the trading task, this can be attributed to the extreme price fluctuations during the COVID-19 event. Both stocks experienced a sharp decline of over 30%, followed by a subsequent rebound. Given STORM’s strong profitability capability, its trading strategy tends to aggressively capture trends, which naturally leads to larger drawdowns and increased volatility as a trade-off for higher potential returns.
@@
Moreover, considering the return and risk simultaneously, STORM performs the best on ASR and CR with an average improvement of 123.94% and 245.45% respectively, over the best baseline result in the PM task, and surpasses the best baseline results across three stocks in the AT task. Such an outstanding ability to balance return and risk is particularly valuable in real-world investment, where achieving high returns while effectively managing risk is crucial for long-term financial success.
Table 4. Factor quality evaluation task results on RankIC and RankICIR (mean ± range, computed across 10 runs).
Models SP500 DJ30
[…]
Future Return Prediction in Portfolio Management Task
LightGBM 0.027 0.274 0.031 0.272
± 0.006 ± 0.084 ± 0.005 ± 0.049
LSTM 0.034 0.333 0.031 0.329
± 0.006 ± 0.042 ± 0.004 ± 0.056
Transformer 0.035 0.340 0.033 0.343
± 0.007 ± 0.078 ± 0.005 ± 0.045
CAFactor 0.037 0.356 0.040 0.380
± 0.005 ± 0.084 ± 0.003 ± 0.043
FactorVAE 0.052 0.543 0.056 0.520
± 0.010 ± 0.122 ± 0.012 ± 0.081
HireVAE 0.057 0.558 0.058 0.563
± 0.006 ± 0.058 ± 0.006 ± 0.053
STORM 0.062 0.673 0.065 0.668
± 0.018 ± 0.155 ± 0.038 ± 0.287
STORM-w/o-TS 0.053 0.513 0.055 0.563
± 0.017 ± 0.145 ± 0.015 ± 0.127
STORM-w/o-CS 0.054 0.522 0.053 0.559
± 0.014 ± 0.118 ± 0.016 ± 0.156
Improvement(%)11footnotemark: 1 8.772 20.609 12.069 18.650
  • 11
    Improvement of STORM over the best-performing baselines.
5.3. Factor Quality (for RQ2)
To demonstrate the effectiveness and quality of the factors, we designed a stock future return prediction task. The goal is not to pursue time-series forecasting accuracy, but to evaluate the predictive quality and robustness of the factors in capturing dynamic market patterns. In this task, STORM shows substantial improvements over six methods across the SP500 and DJ30 datasets. We use rank-based correlation metrics (i.e., RankIC, RankICIR) instead of error-based ones (e.g., MSE, MAE), as they better capture a factor’s ability to preserve the correct cross-sectional ordering of returns. As shown in Table  4, STORM achieves substantial average improvements of 14.690% on SP500 and 15.360% on DJ30, demonstrating its ability to capture more consistent and generalizable signals across diverse market conditions.
Figure 3. (a) Distribution of counts for codebook embedding indices. (b) Hyperparameter experiment results for different codebook sizes.
Factor Diversity. The diversity of factors is essential in financial investment, as it allows the model to capture the multi-dimensional characteristics of the market, avoiding redundancy or over-reliance on a single factor. In the codebook, each embedding represents a cluster center for a category of factor embeddings. To analyze the utilization rate of each factor category, we report the frequency of codebook embedding indices in the testing data. For clarity, we group every five indices into one interval, as illustrated in Figure 3(a). The results indicates a relatively even distribution of factor usage, highlighting the model’s effectiveness to maintain diversity during the factor selection process.
Codebook Design. Furthermore, we conduct a hyperparameter analysis on the DJ30 dataset to examine the effect of different codebook sizes, aiming to balance factor diversity and representational quality. An excessively large codebook may lead to sparse factor representations, limiting the effectiveness of factor utilization, while an overly small codebook could result in insufficient factor differentiation and information redundancy. Following standard practices, we evaluate five different codebook sizes, including 128, 256, 512, 1024, and 2048. As shown in Figure 3(b), a size of 512 achieves the best overall factor prediction performance, offering a favorable trade-off between expressiveness and efficiency.
@@ 5.4. Ablation Study (for RQ3)
We conducted ablation experiments with two simplified versions of STORM, namely STORM-w/o-TS and STORM-w/o-CS, which only extract cross-sectional and time-series factors, respectively. Specifically, the absence of either time-series or cross-sectional factors significantly reduces the model’s effectiveness across various tasks. In the stock return prediction task, STORM outperforms both simplified models, with an average improvement of 16.50% on the RankIC metric and 23.79% on the RankICIR metric, highlighting the enhanced accuracy achieved by incorporating both time-series and cross-sectional factors. In the PM and AT tasks, STORM and two simplified models all performed well. Notably, STORM-w/o-TS exhibited the lowest AVO among the three models, although its profitability did not surpass that of STORM. These findings suggest that methods like FactorVAE (duan2022factorvae), which focus solely on cross-sectional factors, may struggle to capture the full complexity of the financial market. Therefore, integrating both time-series and cross-sectional factors is essential for robust and accurate financial modeling.
@@ 5.5. Discussion on Performance
Efficiency. STORM runs efficiently in real-world daily trading without the ultra-low latency or specialized hardware needed for high-frequency systems, even on standard CPUs. This is enabled by a compact architecture with only a few transformer layers and 15.2M FP32 parameters on DJ30, enabling deployment on edge devices while supporting concurrent market analysis and trading.
Effectiveness. STORM delivers consistently strong performance across diverse tasks, including PM and AT, outperforming baselines that often specialize in one domain but fail to generalize. It balances return and risk, targeting comprehensive investment performance rather than over-optimizing for a single metric. Robustness and versatility across tasks justify the design of our proposed model.
Generalization. STORM generalizes across markets of varying scales and characteristics, excelling on both DJ30 (large-cap industrial blue chips) and SP500 (broad, diversified industries). These representative U.S. datasets cover concentrated and diversified indices, and their depth makes them a strong proxy for other markets like China or the U.K.. The flexible encoder-decoder architecture also allows using additional layers like Mamba  (gu2023mamba; mamba2) and ensures applicability to diverse trading environments and asset classes.
@@ 6. Conclusions and Future Work
In this paper, we propose STORM, a spatio-temporal latent factor model leveraging a dual VQ-VAE framework to learn both time-series and cross-sectional factors. STORM ensures factor orthogonality and diversity, crucial for effective factor selection in financial trading. Unlike traditional factors, STORM learns high-dimensional latent vectors that, though less interpretable, capture richer market structures. Extensive experiments demonstrate STORM’s superiority over the state-of-the-art methods across a prediction task and two downstream tasks, showcasing robust performance and effectiveness. In the future, we plan to extend STORM by incorporating side information like news, to explore the effects of exogenous factors.
Acknowledgments
This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL). It is also supported, in part, by Jinan-NTU Green Technology Research Institute (GreenTRI).